# Krylov exponents and power spectra for maximal quantum chaos: an EFT approach

### v2

@@ 1 Introduction and motivation
with $N\gg 1$ a measure of the number of degrees of freedom, and $a$ a theory-dependent order $1$ constant. The OTOC decays until the scrambling time $t_{s}\sim\lambda_{L}^{-1}\log N$, and the decay rate is controlled by the Lyapunov exponent $\lambda_{L}$. The Lyapunov exponent is upper-bounded by the Maldacena-Shenker-Stanford (MSS) bound Maldacena:2015waa  [^2] We work in natural units, $k_{B}=\hbar=1.$
@@
Systems that saturate this bound are referred to as maximally chaotic. Examples include holographic models, such as the Sachdev-Ye-Kitaev (SYK) model in the large $N$, low temperature limit Shenker:2013pqa ; Roberts:2014ifa ; Maldacena:2016hyu .
@@
To capture this universal feature of chaotic dynamics in these systems, an effective field theory (EFT) approach has been developed in Blake:2017ris ; Blake:2021wqj . The key assumption is that the growth of generic few-body operators comes from their coupling to hydrodynamic modes. The construction relies on building an action for these hydrodynamic modes to all orders in a derivative expansion, referred to as quantum hydrodynamics, which is based on symmetry principles following the methods outlined in Crossley:2015evo ; Glorioso:2017fpd . In addition to the symmetries needed for quantum hydrodynamics, a crucial structural ingredient in this EFT construction is a postulated shift symmetry, which acts on the hydrodynamic variables and constrains the allowed interactions. This shift symmetry was shown to ensure the growth of the OTOCs in $0+1$ dimensional systems Blake:2017ris . Note that while this EFT successfully captures key features of maximally chaotic dynamics as probed by the OTOC, it does not, for instance, describe the full microscopic structure of operator growth, nor does it generalise straightforwardly to dynamics beyond leading order in $1/N$.
@@
Complementing OTOCs, the Universal Operator Growth Hypothesis (UOGH) and the associated Krylov complexity have emerged as diagnostics of operator growth and quantum chaos. See Parker:2018yvk for the original proposal, Nandy:2025ktk ; Baiguera:2025dkc ; Rabinovici:2025otw for comprehensive reviews, and Caputa:2021sib ; Rabinovici:2022beu ; Erdmenger:2023wjg ; Hashimoto:2023swv ; Camargo:2023eev ; Camargo:2024deu ; Alishahiha:2024vbf ; Craps:2024suj ; Malvimat:2024vhr ; Chattopadhyay:2024pdj ; Jian:2020qpp ; Hornedal:2022pkc ; He:2022ryk ; Bhattacharjee:2022vlt ; Baggioli:2024wbz ; Chen:2024imd ; Ambrosini:2024sre ; Caputa:2024sux for recent work exploring their connection to chaos. Krylov complexity quantifies how an operator spreads in Krylov space, the space spanned by nested commutators of the Hamiltonian with an initial operator, constructed via the Lanczos algorithm. The resulting Lanczos coefficients $b_{n}$ define a tridiagonal representation of the Liouvillian, and in chaotic systems in the thermodynamic limit are conjectured to grow linearly, $b_{n}\sim\alpha n$ as $n\to\infty$. This linear growth yields, under certain conditions, exponential growth of the Krylov complexity, $K(t)\propto\exp(\lambda_{K}t)$, governed by the Krylov exponent $\lambda_{K}$. The latter is conjectured to upper-bound the Lyapunov exponent derived from the OTOC.
More specifically, at finite temperature, the Krylov exponent has been conjectured, and proven at infinite temperature, to provide a tighter bound on chaos than the MSS bound
@@
Supporting evidence for this conjecture was given in Avdoshkin:2019trj ; Gu:2021xaj . However, the bound also highlights a limitation of Krylov complexity as a diagnostic for chaos. In particular, systems with maximal $\lambda_{K}$ and submaximal $\lambda_{L}$ have been constructed Chapman:2024pdw ; Bhattacharjee:2022vlt , indicating that Krylov complexity can, in some cases, fail to detect the submaximal and even integrable nature of the model.
More issues with this bound come from quantum field theory (QFT) considerations. Specifically, correlation functions of generic operators in QFTs exhibit short-distance singularities when the operator insertions approach the same spacetime point, known as the coincident point limit. For a smooth Lanczos sequence, such singularities in the two-point function fully determine the asymptotes of the Lanczos coefficients. For example, in Dymarsky:2021bjq , maximal linear growth of the Lanczos sequence was obtained for integrable and free field theories, indicating that this behaviour should not always be associated with the presence of chaos. However, by introducing an energy cutoff, e.g. a UV cutoff, the value of $\lambda_{K}$ can distinguish integrable and chaotic field theories Camargo:2022rnt ; Avdoshkin:2022xuw .
@@
Our strategy for testing this bound is to find two-point functions that satisfy the shift symmetry constraints, as well as a minimal set of physical requirements: KMS symmetry, reality, and unitarity. For these solutions, we then compute the corresponding Lanczos coefficients. In the case where these coefficients are a smooth function of $n$, when extended from the integers to the reals, the asymptotic growth of the found $b_{n}$ determines the associated Krylov exponent $\lambda_{K}$ Parker:2018yvk . A simple example that solves the shift symmetry constraints is the conformal correlator $g_{\hat{V}}(t)=\sinh^{-2\Delta_{V}}(\pi t/\beta)$. Its Lanczos coefficients are known: they are asymptotically linear, $b_{n}\sim\frac{\pi}{\beta}n$, giving $\lambda_{K}=2\pi T=\lambda_{L}$ and thus this example satisfies the UOGH and the exponent bound eq. (3). However, not all solutions of the shift symmetry constraints that we construct satisfy the exponent bound. In particular, we find that, in the considered examples, the Krylov exponent $\lambda_{K}$ takes one of two values: $2\pi T$ or $\pi T$. The latter case, $\lambda_{K}=\pi T$, appears to be in tension with eq. (3), as well as with the argument for the bound given in Gu:2021xaj for a restricted class of models. However, as we will discuss in subsection 3.5, the argument of Gu:2021xaj relies on a specific factorisation structure of the OTOC, which is not assumed in the EFT construction. Our result raises the possibility that the conjecture may be more sensitive to dynamical details than previously appreciated.
This paper is organised as follows. In section 2, we start by reviewing the effective field theory (EFT) description of chaos proposed in Blake:2017ris ; Blake:2021wqj . We also include a summary of the UOGH and the definition of the Krylov exponent. Building on this, we summarise the various constraints that are to be imposed on two-point functions of generic operators that arise in this framework. This includes the main ingredient of the EFT construction, a constraint on two-point functions due to the presence of a shift symmetry. In section 3, we solve the shift symmetry constraints for several illustrative examples. Computing the corresponding Krylov exponent, we encounter cases that exhibit both maximal and submaximal growth. Finally, we provide a critical assessment of these results, contrasting the defining assumptions of the EFT and Krylov complexity. In section 4, we turn to a general analysis of the solutions of the shift symmetry constraints and derive a thermal product-like formula for the power spectrum for a set of examples. We show that the solution space of the shift symmetry constraints can be identified with a subclass of Meijer G-functions. Finally, we complement the analysis in the main text with several technical appendices.
@@ 2 Operator growth in the EFT for many-body maximal quantum chaos
In this section, we review both the effective field theory approach to maximally chaotic many-body quantum systems, as proposed by Blake:2017ris ; Blake:2021wqj , and the UOGH and Krylov complexity. We also explain the constraints that are to be imposed on two-point functions of generic operators in the EFT.
@@ 2.1 The hydrodynamical EFT of maximal chaos
In Blake:2017ris ; Blake:2021wqj , the authors developed a hydrodynamic effective theory for maximally chaotic quantum systems, capturing the universal features of operator scrambling in large-$N$ many-body systems. Let us start by briefly reviewing the basic ingredients of this formalism.
@@
The most general action that can come from such a procedure has to be consistent with the symmetries and physical constraints of the system. Some of these constraints, such as those associated with conserved quantities, are model-dependent. Others are universal and follow from fundamental principles like unitarity and CPT invariance of the underlying quantum theory. Moreover, if the initial state is thermal, $\rho_{\beta}=Z(\beta)^{-1}e^{-\beta{H}}$, where H{H} is the Hamiltonian, the effective action also satisfies the Kubo-Martin-Schwinger (KMS) symmetry. See Crossley:2015evo ; Glorioso:2017fpd for more details on this construction and Liu:2018kfw for a review. The explicit expression for the most general action consistent with these requirements is rather intricate and does not play a direct role in the arguments presented in this paper. We therefore refer the reader to Blake:2017ris for further details.
@@
A crucial input in Blake:2017ris ; Blake:2021wqj to render the effective action maximally chaotic is requiring that the effective action of the hydrodynamic mode is invariant under a shift symmetry
@@
The framework of Blake:2017ris ; Blake:2021wqj further builds on the intuition that the evolution of a generic few-body bosonic operator $V(t)$ at finite temperature can be decomposed into a “bare” component $\hat{V}(t)$, which captures its microscopic degrees of freedom, and a hydrodynamic “cloud” arising from coupling to a hydro mode $\epsilon$ associated with energy conservation
@@
with coefficients $c_{nm}$ encoding the effective low-energy dynamics. This decomposition is depicted in the left panel of Figure 1. Crucially, Blake:2017ris ; Blake:2021wqj assume that bare operators $\hat{V}$, $\hat{W}$ that involve different degrees of freedom can only communicate with themselves: $\langle\hat{V}(t_{1})\hat{W}(t_{2})\rangle_{\beta}=0$ for $V\neq W$.
@@
Turning to four-point functions, the key result of Blake:2017ris ; Blake:2021wqj is that at leading order $\epsilon(t)$, the OTOCs can be expressed in terms of effective vertices coupling two bare operators to the hydrodynamic $B_{\hat{V}}$
@@
As shown in Blake:2017ris ; Blake:2021wqj , the shift symmetry together with the constraints in eq. (13) are the essential ingredients in the effective theory construction that ensure exponential growth appears only in OTOCs, and not in time-ordered correlators (TOCs). In TOCs, the symmetry enforces a cancellation of the growing contributions, so they do not exhibit any exponential sensitivity to time separations. In contrast, the non-trivial time ordering in OTOCs obstructs this cancellation, leading to exponential growth in the connected piece of the OTOC, as in eq. (1). In this way, the hydrodynamic effective theory of Blake:2017ris ; Blake:2021wqj reproduces the universal hallmark of maximal quantum chaos: OTOCs grow exponentially in time with $\lambda_{L}=2\pi T$, while TOCs do not grow exponentially. In this paper, unless otherwise stated, $\lambda_{L}$ always denotes its value in the EFT: $\lambda_{L}=2\pi T.$
@@ 2.2 The Universal Operator Growth Hypothesis and Krylov exponents
A proposal to quantify operator growth was put forward in Parker:2018yvk , where the idea is to study the evolution of $\hat{V}(t)$ by constructing a Krylov basis from the nested commutators $\{\mathrm{ad}_{H}^{n}\hat{V}\}_{n}$. Using the Lanczos algorithm, a variation of the Gram-Schmidt orthogonalisation specialised to Hermitian dynamics, one obtains an orthonormal basis $\{|\hat{V}_{n})\}_{n}$ and a sequence of real positive Lanczos coefficients $\{b_{n}\}$. In this basis, the dynamics of the operator are equivalent to a tight-binding chain with variable hopping
@@
The central conjecture of Parker:2018yvk , the UOGH, is that, in quantum chaotic many-body systems, the Lanczos coefficients grow as fast as possible. Specifically, the coefficients grow linearly in $n$ at asymptotically large values of $n$
@@
The authors in Parker:2018yvk conjectured that the Krylov exponent provides a bound on chaos, proposing that the maximal rate of operator growth bounds the maximal rate of scrambling as measured by the OTOC
@@
If the Lanczos coefficients asymptote to $b_{n}\sim\alpha n$ as $n\to\infty$, and when the associated Krylov complexity is exponentially growting, we can identify $2\alpha=\lambda_{K}$. If not, the Krylov exponent is obtained by solving (16) and computing the Krylov complexity directly. At infinite temperature, this conjectured relation between $\alpha$ and the Lyapunov exponent has been proven Parker:2018yvk . At finite temperature, the Lanczos coefficients, and hence their slope $\alpha$, acquire a temperature dependence. In Parker:2018yvk , by adopting the Wightman inner product, the authors conjecture that the bound in eq. (19) still applies and is verified in several examples, including Schwarzian quantum mechanics and the large-$N$, large-$q$ limit of SYK (taking $N$ large first). In particular, in the former, the Krylov exponent was shown to saturate the bound: $\lambda_{L}=\lambda_{K}$.
@@
the Lanczos coefficients display an asymptotic behaviour of the form $b_{n}^{2}\sim n^{\rho}$ lubinsky1993update , where the proportionality constant is related to the decay rate $\omega_{0}$ in eq. (22). For linear growth ($\rho=2$), the asymptotic growth rate of the Lanczos coefficients $b_{n}$ in eq. (17) and exponential decay rate of the power spectrum $\Phi(\omega)$ are related as $\omega_{0}=2\alpha/\pi$.
@@
In particular, the analyticity properties of $C(t)$ are related to the properties of the other corners of the triangle, most notably; linear growth of $b_{n}$ implies that $C(t)$ is analytic in a strip of width $1/\omega_{0}$ in the complex time plane, with $\omega_{0}=2\alpha/\pi$ Parker:2018yvk . It is important to emphasise that the relation $\pi\omega_{0}=2\alpha=\lambda_{K}$ only holds when $b_{n}$ is asymptotically linear and when the associated Krylov complexity grows exponentially. When the Lanczos coefficients $b_{n}$ are not smooth functions of $n$, when $n$ is extended from the integers to the reals, the simple correspondence may no longer hold, and this was explored in Dymarsky:2021bjq ; Camargo:2022rnt ; Avdoshkin:2022xuw ; Anegawa:2024wov ; Anegawa:2024yia . Persistent staggering, defined in Avdoshkin:2022xuw as when $b_{n}$ is asymptotically linear but with different intercepts for odd and even $n$, is an example of non-smooth behaviour that has been studied.
@@
At finite temperature, using the relation between the slope of the Lanczos coefficients and the high-frequency decay of the spectral function, Parker:2018yvk showed that the decay rate $\alpha$ satisfies the inequality
@@
The applicability of this bound relies on the analytic properties of the two-point function and, crucially, on the behaviour of the Lanczos coefficients. In particular, it assumes that the coefficients $b_{n}$ are asymptotically linear in $n$ with slope $\alpha$, in which case $2\alpha=\lambda_{K}=\pi\omega_{0}$ Parker:2018yvk . However, when the Lanczos coefficients are not smooth, such as when they display staggering, this correspondence breaks down. In such cases, $\lambda_{K}$ can still be defined directly from the growth of Krylov complexity $K(t)$. This motivates a refined version of the quantum chaos bound that generalises eq. (24) and remains applicable even when $b_{n}$ is non-smooth:
@@
Further support for this type of bound, under certain technical assumptions, was later provided by Gu:2021xaj , who proposed $\lambda_{L}\leq\pi\omega_{0}\leq 2\pi T$, see also Avdoshkin:2019trj .

### v3

@@ 1 Introduction and motivation
with $N\gg 1$ a measure of the number of degrees of freedom, and $a$ a theory-dependent order $1$ constant. The OTOC decays until the scrambling time $t_{s}\sim\lambda_{L}^{-1}\log N$, and the decay rate is controlled by the Lyapunov exponent $\lambda_{L}$. The Lyapunov exponent is upper-bounded by the Maldacena-Shenker-Stanford (MSS) bound Maldacena et al. (2016) [^2] We work in natural units, $k_{B}=\hbar=1.$
@@
Systems that saturate this bound are referred to as maximally chaotic. Examples include holographic models, such as the Sachdev-Ye-Kitaev (SYK) model in the large $N$, low temperature limit Shenker and Stanford (2014); Roberts and Stanford (2015); Maldacena and Stanford (2016).
@@
To capture this universal feature of chaotic dynamics in these systems, an effective field theory (EFT) approach has been developed in Blake et al. (2018); Blake and Liu (2021). The key assumption is that the growth of generic few-body operators comes from their coupling to hydrodynamic modes. The construction relies on building an action for these hydrodynamic modes to all orders in a derivative expansion, referred to as quantum hydrodynamics, which is based on symmetry principles following the methods outlined in Crossley et al. (2017); Glorioso et al. (2017). In addition to the symmetries needed for quantum hydrodynamics, a crucial structural ingredient in this EFT construction is a postulated shift symmetry, which acts on the hydrodynamic variables and constrains the allowed interactions. This shift symmetry was shown to ensure the growth of the OTOCs in $0+1$ dimensional systems Blake et al. (2018). Note that while this EFT successfully captures key features of maximally chaotic dynamics as probed by the OTOC, it does not, for instance, describe the full microscopic structure of operator growth, nor does it generalise straightforwardly to dynamics beyond leading order in $1/N$.
@@
Complementing OTOCs, the Universal Operator Growth Hypothesis (UOGH) and the associated Krylov complexity have emerged as diagnostics of operator growth and quantum chaos. See Parker et al. (2019) for the original proposal, Nandy et al. (2025); Baiguera et al. (2025); Rabinovici et al. (2025) for comprehensive reviews, and Caputa et al. (2022); Rabinovici et al. (2022); Erdmenger et al. (2023); Hashimoto et al. (2023); Camargo et al. (2024b, a); Alishahiha et al. (2024); Craps et al. (2025); Malvimat et al. (2025); Chattopadhyay et al. (2024); Jian et al. (2021); Hörnedal et al. (2022); He et al. (2022); Bhattacharjee et al. (2022); Baggioli et al. (2025); Chen et al. (2025); Ambrosini et al. (2024); Caputa et al. (2024) for recent work exploring their connection to chaos. Krylov complexity quantifies how an operator spreads in Krylov space, the space spanned by nested commutators of the Hamiltonian with an initial operator, constructed via the Lanczos algorithm. The resulting Lanczos coefficients $b_{n}$ define a tridiagonal representation of the Liouvillian, and in chaotic systems in the thermodynamic limit are conjectured to grow linearly, $b_{n}\sim\alpha n$ as $n\to\infty$. This linear growth yields, under certain conditions, exponential growth of the Krylov complexity, $K(t)\propto\exp(\lambda_{K}t)$, governed by the Krylov exponent $\lambda_{K}$. The latter is conjectured to upper-bound the Lyapunov exponent derived from the OTOC.
More specifically, at finite temperature, the Krylov exponent has been conjectured, and proven at infinite temperature, to provide a tighter bound on chaos than the MSS bound [^3] Note that at finite temperature, this statement depends on the chosen regularization scheme used to define the thermal OTOC. Therefore, the Lyapunov exponent $\lambda_{L}$ can in general depend on this choice, see e.g. Romero-Bermúdez et al. (2019). We will comment on this when reviewing the conjecture in detail around eq. (19).
@@
Supporting evidence for this conjecture was given in Avdoshkin and Dymarsky (2020); Gu et al. (2022). However, the bound also highlights a limitation of Krylov complexity as a diagnostic for chaos. In particular, systems with maximal $\lambda_{K}$ and submaximal $\lambda_{L}$ have been constructed Chapman et al. (2025); Bhattacharjee et al. (2022), indicating that Krylov complexity can, in some cases, fail to detect the submaximal and even integrable nature of the model.
More issues with this bound come from quantum field theory (QFT) considerations. Specifically, correlation functions of generic operators in QFTs exhibit short-distance singularities when the operator insertions approach the same spacetime point, known as the coincident point limit. For a smooth Lanczos sequence, such singularities in the two-point function fully determine the asymptotes of the Lanczos coefficients. For example, in Dymarsky and Smolkin (2021), maximal linear growth of the Lanczos sequence was obtained for integrable and free field theories, indicating that this behaviour should not always be associated with the presence of chaos. However, by introducing an energy cutoff, e.g. a UV cutoff, the value of $\lambda_{K}$ can distinguish integrable and chaotic field theories Camargo et al. (2023); Avdoshkin et al. (2022).
@@
Our strategy for examining whether maximal chaos as seen through the OTOC implies maximal Krylov growth in the EFT is to find two-point functions that satisfy the shift symmetry constraints, as well as a minimal set of physical requirements: KMS symmetry, reality, and unitarity. For these solutions, we then compute the corresponding Lanczos coefficients. In the case where these coefficients are a smooth function of $n$, when extended from the integers to the reals, the asymptotic growth of the found $b_{n}$ determines the associated Krylov exponent $\lambda_{K}$ Parker et al. (2019). A simple example that solves the shift symmetry constraints is the conformal correlator $g_{\hat{V}}(t)=\sinh^{-2\Delta_{V}}(\pi t/\beta)$. Its Lanczos coefficients are known: they are asymptotically linear, $b_{n}\sim\frac{\pi}{\beta}n$, giving $\lambda_{K}=2\pi T=\lambda_{L}$ and thus this example satisfies the UOGH and the exponent bound eq. (3). However, not all solutions of the shift symmetry constraints that we construct satisfy the exponent bound. In particular, we find that, in the considered examples, the Krylov exponent $\lambda_{K}$ takes one of two values: $2\pi T$ or $\pi T$. The latter case, $\lambda_{K}=\pi T$, suggests that the shift symmetry is not enough to ensure maximal Krylov growth within our working assumptions. Moreover, we will comment on the possible tension with eq. (3), as well as with the argument for the bound given in Gu et al. (2022) for a restricted class of models. However, as we will discuss in subsection 3.5, the argument of Gu et al. (2022) relies on a specific factorisation structure of the OTOC, which is not assumed in the EFT construction. Our result raises the possibility that the conjecture may be more sensitive to dynamical details than previously appreciated.
This paper is organised as follows. In section 2, we start by reviewing the effective field theory (EFT) description of chaos proposed in Blake et al. (2018); Blake and Liu (2021). We also include a summary of the UOGH and the definition of the Krylov exponent. Building on this, we summarise the various constraints that are to be imposed on two-point functions of generic operators that arise in this framework. This includes the main ingredient of the EFT construction, a constraint on two-point functions due to the presence of a shift symmetry. In section 3, we solve the shift symmetry constraints for several illustrative examples. Computing the corresponding Krylov exponent, we encounter cases that exhibit both maximal and submaximal growth. Finally, we provide a critical assessment of these results, contrasting the defining assumptions of the EFT and Krylov complexity. In section 4, we turn to a general analysis of the solutions of the shift symmetry constraints and derive a thermal product-like formula for the power spectrum for a set of examples. We show that the solution space of the shift symmetry constraints can be identified with a subclass of Meijer G-functions. Finally, we complement the analysis in the main text with several technical appendices.
@@ 2 Operator growth in the EFT for many-body maximal quantum chaos
In this section, we review both the effective field theory approach to maximally chaotic many-body quantum systems, as proposed by Blake et al. (2018); Blake and Liu (2021), and the UOGH and Krylov complexity. We also explain the constraints that are to be imposed on two-point functions of generic operators in the EFT.
@@ 2.1 The hydrodynamical EFT of maximal chaos
In Blake et al. (2018); Blake and Liu (2021), the authors developed a hydrodynamic effective theory for maximally chaotic quantum systems, capturing the universal features of operator scrambling in large-$N$ many-body systems. Let us start by briefly reviewing the basic ingredients of this formalism.
@@
The most general action that can come from such a procedure has to be consistent with the symmetries and physical constraints of the system. Some of these constraints, such as those associated with conserved quantities, are model-dependent. Others are universal and follow from fundamental principles like unitarity and CPT invariance of the underlying quantum theory. Moreover, if the initial state is thermal, $\rho_{\beta}=Z(\beta)^{-1}e^{-\beta{H}}$, where H{H} is the Hamiltonian, the effective action also satisfies the Kubo-Martin-Schwinger (KMS) symmetry. See Crossley et al. (2017); Glorioso et al. (2017) for more details on this construction and Liu and Glorioso (2018) for a review. The explicit expression for the most general action consistent with these requirements is rather intricate and does not play a direct role in the arguments presented in this paper. We therefore refer the reader to Blake et al. (2018) for further details.
@@
A crucial input in Blake et al. (2018); Blake and Liu (2021) to render the effective action maximally chaotic is requiring that the effective action of the hydrodynamic mode is invariant under a shift symmetry
@@
The framework of Blake et al. (2018); Blake and Liu (2021) further builds on the intuition that the evolution of a generic few-body bosonic operator $V(t)$ at finite temperature can be decomposed into a “bare” component $\hat{V}(t)$, which captures its microscopic degrees of freedom, and a hydrodynamic “cloud” arising from coupling to a hydro mode $\epsilon$ associated with energy conservation
@@
with coefficients $c_{nm}$ encoding the effective low-energy dynamics. This decomposition is depicted in the left panel of Figure 1. Crucially, Blake et al. (2018); Blake and Liu (2021) assume that bare operators $\hat{V}$, $\hat{W}$ that involve different degrees of freedom can only communicate with themselves: $\langle\hat{V}(t_{1})\hat{W}(t_{2})\rangle_{\beta}=0$ for $V\neq W$.
@@
Turning to four-point functions, the key result of Blake et al. (2018); Blake and Liu (2021) is that at leading order $\epsilon(t)$, the OTOCs can be expressed in terms of effective vertices coupling two bare operators to the hydrodynamic $B_{\hat{V}}$
@@
As shown in Blake et al. (2018); Blake and Liu (2021), the shift symmetry together with the constraints in eq. (13) are the essential ingredients in the effective theory construction that ensure exponential growth appears only in OTOCs, and not in time-ordered correlators (TOCs). In TOCs, the symmetry enforces a cancellation of the growing contributions, so they do not exhibit any exponential sensitivity to time separations. In contrast, the non-trivial time ordering in OTOCs obstructs this cancellation, leading to exponential growth in the connected piece of the OTOC, as in eq. (1). In this way, the hydrodynamic effective theory of Blake et al. (2018); Blake and Liu (2021) reproduces the universal hallmark of maximal quantum chaos: OTOCs grow exponentially in time with $\lambda_{L}=2\pi T$, while TOCs do not grow exponentially. In this paper, unless otherwise stated, $\lambda_{L}$ always denotes its value in the EFT: $\lambda_{L}=2\pi T.$
@@ 2.2 The Universal Operator Growth Hypothesis and Krylov exponents
A proposal to quantify operator growth was put forward in Parker et al. (2019), where the idea is to study the evolution of $\hat{V}(t)$ by constructing a Krylov basis from the nested commutators $\{\mathrm{ad}_{H}^{n}\hat{V}\}_{n}$. Using the Lanczos algorithm, a variation of the Gram-Schmidt orthogonalisation specialised to Hermitian dynamics, one obtains an orthonormal basis $\{|\hat{V}_{n})\}_{n}$ and a sequence of real positive Lanczos coefficients $\{b_{n}\}$. In this basis, the dynamics of the operator are equivalent to a tight-binding chain with variable hopping
@@
The central conjecture of Parker et al. (2019), the UOGH, is that, in quantum chaotic many-body systems, the Lanczos coefficients grow as fast as possible. Specifically, the coefficients grow linearly in $n$ at asymptotically large values of $n$
@@
The authors in Parker et al. (2019) conjectured that the Krylov exponent provides a bound on chaos, proposing that the maximal rate of operator growth bounds the maximal rate of scrambling as measured by the OTOC [^4] The original conjecture as stated in Parker et al. (2019) considers the thermal OTOCs as defined in Maldacena et al. (2016) which are of the form $\operatorname{Tr}\!\left[y\,V(0)\,y\,W(t)\,y\,V(0)\,y\,W(t)\right]$ with $y^{4}=\rho_{\beta}$. In general, the Lyapunov spectrum extracted from the thermal OTOCs can depend on the choice of regularization scheme. An example where such contour dependence was considered in weakly coupled field theories is Romero-Bermúdez et al. (2019). On the other hand, in large-$N$ models such as SYK, the Lyapunov spectrum does not exhibit contour dependence Shenker and Stanford (2015); Chowdhury and Swingle (2017); Romero-Bermúdez et al. (2019), consistent with the expectation that in maximally chaotic systems the exponent is universal. In this work, we will extract $\lambda_{L}$ from the thermal OTOC used in Blake et al. (2018), which of the form $\text{Tr}\!\left[\rho_{\beta}\,V(0)\,W(t)\,V(0)\,W(t)\right]$, and interpret comparisons to $\lambda_{K}$ accordingly.
@@
If the Lanczos coefficients asymptote to $b_{n}\sim\alpha n$ as $n\to\infty$, and when the associated Krylov complexity is exponentially growting, we can identify $2\alpha=\lambda_{K}$. If not, the Krylov exponent is obtained by solving (16) and computing the Krylov complexity directly. At infinite temperature, this conjectured relation between $\alpha$ and the Lyapunov exponent has been proven Parker et al. (2019). At finite temperature, the Lanczos coefficients, and hence their slope $\alpha$, acquire a temperature dependence. In Parker et al. (2019), by adopting the Wightman inner product, the authors conjecture that the bound in eq. (19) still applies and is verified in several examples, including Schwarzian quantum mechanics and the large-$N$, large-$q$ limit of SYK (taking $N$ large first). In particular, in the former, the Krylov exponent was shown to saturate the bound: $\lambda_{L}=\lambda_{K}$.
@@
the Lanczos coefficients display an asymptotic behaviour of the form $b_{n}^{2}\sim n^{\rho}$ Lubinsky (1993), where the proportionality constant is related to the decay rate $\omega_{0}$ in eq. (22). For linear growth ($\rho=2$), the asymptotic growth rate of the Lanczos coefficients $b_{n}$ in eq. (17) and exponential decay rate of the power spectrum $\Phi(\omega)$ are related as $\omega_{0}=2\alpha/\pi$.
@@
In particular, the analyticity properties of $C(t)$ are related to the properties of the other corners of the triangle, most notably; linear growth of $b_{n}$ implies that $C(t)$ is analytic in a strip of width $1/\omega_{0}$ in the complex time plane, with $\omega_{0}=2\alpha/\pi$ Parker et al. (2019). It is important to emphasise that the relation $\pi\omega_{0}=2\alpha=\lambda_{K}$ only holds when $b_{n}$ is asymptotically linear and when the associated Krylov complexity grows exponentially. When the Lanczos coefficients $b_{n}$ are not smooth functions of $n$, when $n$ is extended from the integers to the reals, the simple correspondence may no longer hold, and this was explored in Dymarsky and Smolkin (2021); Camargo et al. (2023); Avdoshkin et al. (2022); Anegawa et al. (2024); Anegawa and Watanabe (2024). Persistent staggering, defined in Avdoshkin et al. (2022) as when $b_{n}$ is asymptotically linear but with different intercepts for odd and even $n$, is an example of non-smooth behaviour that has been studied.
@@
At finite temperature, using the relation between the slope of the Lanczos coefficients and the high-frequency decay of the spectral function, Parker et al. (2019) showed that the decay rate $\alpha$ satisfies the inequality
@@
The applicability of this bound relies on the analytic properties of the two-point function and, crucially, on the behaviour of the Lanczos coefficients. In particular, it assumes that the coefficients $b_{n}$ are asymptotically linear in $n$ with slope $\alpha$, in which case $2\alpha=\lambda_{K}=\pi\omega_{0}$ Parker et al. (2019). However, when the Lanczos coefficients are not smooth, such as when they display staggering, this correspondence breaks down. In such cases, $\lambda_{K}$ can still be defined directly from the growth of Krylov complexity $K(t)$. This motivates a refined version of the quantum chaos bound that generalises eq. (24) and remains applicable even when $b_{n}$ is non-smooth:
@@
Let us stress that this inequality should be interpreted with care, as highlighted in footnote 4. Further support for this type of bound, under certain technical assumptions, was later provided by Gu et al. (2022), who proposed $\lambda_{L}\leq\pi\omega_{0}\leq 2\pi T$, see also Avdoshkin and Dymarsky (2020).